OpenStudy (anonymous):

can someone help me solve the equation and check if it has an extraneous solution?

OpenStudy (anonymous):

I can help

OpenStudy (anonymous):

2√x+5+7=11

OpenStudy (anonymous):

That is extraneous but I think that this would be easier to solve:

OpenStudy (anonymous):

2√x + 4 + 10= 6

OpenStudy (anonymous):

I'll send you the steps

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

2√x + 4 + 10= 6 Subtract both sides by 10 2√x + 4= -4 Divide both sides by 2 √x + 4= -2 Square both sides x + 4= 4 Subtract both sides by 4 x= 0

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Pls medal

OpenStudy (anonymous):

ill help u with task 3

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

part 1

OpenStudy (anonymous):

i dont understand part 3

OpenStudy (anonymous):

put a nice round number on the first blank

OpenStudy (anonymous):

(5,10,20) or any like that

OpenStudy (anonymous):

or even 1 if you want to be lazy

OpenStudy (anonymous):

for the 2nd blank put another nice number

OpenStudy (anonymous):

@NiniNunu

OpenStudy (anonymous):

???